A water spray system discharge water to the atmosphere. Water is released to the atmosphere by two identical nozzles at the same velocity. The pipe diameter is 6 inch and the nozzle diameter are 3 inches. If the vertical reaction force on the flange is 14 kN in magnitude pointing down, determine the velocity V1 within the pipe in m/s. Neglect gravitational force and inlet pressure is P1=250kPa gage.

3400 4000 Sprag Head Nozzle Norrte flanse I TK Ville I pipe pipe
Answer #1

m = 8x4 * (6*0,0254)? * in = 18.24 #V, m/sec D=6inches D=6* 0.0254 m ㅍ putting it in en-eq - lin 40 18:24 * V *[ vsmo #M] =1